THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 92 — 30 March 2008

http://www.astro.keele.ac.uk/MCnews

Editors: Jacco van Loon and Snežana Stanimirović

Editorial

Dear Colleagues,

It is our pleasure to present you the 92^{nd} issue of the Magellanic Clouds Newsletter. It is an exciting one, with two papers on SNR 0509−675 and several papers on evolved massive stars including more results from the VLT-FLAMES survey as well as some high-energy studies. Pulsating stars used as distance indicators keep puzzling us, and several papers have appeared in the recent literature on this topic. The origin and properties of the Magellanic Stream, Bridge and Clouds themselves are also receiving plenty of well-deserved attention.

All the more reason to come to IAU Symposium 256 — http://www.astro.keele.ac.uk/iaus256. Please have your abstracts submitted by 31 March to receive full consideration. The registration and accommodation booking form is expected to go live within the next few days. If you have preregistered you will be notified when this happens, and soon receive the second announcement.

There is a great opportunity to work as a postdoc on the Spitzer spectroscopic survey of the LMC, in Manchester. Please see the advertisement at the end of this Newsletter for more details.

The next issue will be distributed on the 1^{st} of June 2008; the deadline for contributions is the 31^{st} of May.

Editorially Yours,

Jacco van Loon and Snežana Stanimirović

Refereed Journal Papers

Spectral Identification of an Ancient Supernova using Light Echoes in the LMC

A. Rest^{1,2,3}, T. Matheson^{4}, S. Blondin^{5}, M. Bergmann^{6}, D. L. Welch^{7}, N. B. Suntzeff^{8}, R. C. Smith^{1}, K. Olsen^{1}, J.

L. Prieto^{9}, A. Garg^{2}, P. Challis^{5}, C. Stubbs^{2,5}, M. Hicken^{5}, M. Modjaz^{5}, W. M. Wood-Vasey^{5}, A. Zenteno^{1}, G. Damke^{1}, A. Newman^{10}, M. Huber^{11}, K. H. Cook^{11}, S. Nikolaev^{11}, A. C. Becker^{12}, A. Miceli^{12}, R. Covarrubias^{12,13},

L. Morelli^{14}, G. Pignata^{14,15}, A. Clocchiatti^{14}, D. Minniti^{14} and R. J. Foley^{16}

1Cerro Tololo Inter-American Observatory, Chile 2Physics Department, Harvard University, USA 3Goldberg Fellow 4National Optical Astronomy Observatory, USA 5Harvard-Smithsonian Center for Astrophysics, USA 6Gemini Observatory 7McMaster University, Canada 8Texas A&M University, USA 9Ohio State University, USA 10Washington University, USA 11Lawrence Livermore National Laboratory, USA 12University of Washington, USA 13Las Campanas Observatory, Chile 14Pontificia Universidad Catolica de Chile, Chile 15Universidad de Chile, Chile 16University of California, Berkeley, USA

We report the successful identification of the type of the supernova responsible for the supernova remnant SNR 0509−675 in the Large Magellanic Cloud (LMC) using Gemini spectra of surrounding light echoes. The ability to classify outbursts associated with centuries-old remnants provides a new window into several aspects of supernova research and is likely to be successful in providing new constraints on additional LMC supernovae as well as their historical counterparts in the Milky Way Galaxy (MWG). The combined spectrum of echo light from SNR 0509−675 shows broad emission and absorption lines consistent with a supernova (SN) spectrum. We create a spectral library consisting of 26 SNe Ia and 6 SN Ib/c that are time-integrated, dust-scattered by LMC dust, and reddened by the LMC and MWG. We fit these SN templates to the observed light echo spectrum using χ^{2} minimization as well as correlation techniques, and we find that overluminous 91T-like SNe Ia with Δm_{15} < 0.9 match the observed spectrum best.

Accepted for publication in ApJ Available from arXiv:0801.4762

Variations of the mid-IR aromatic features inside and among galaxies

Frédéric Galliano^{1,2}, Suzanne C. Madden^{3}, Alexander G. G. M. Tielens^{4}, Els Peeters^{5,6,7} and Anthony P. Jones^{8}

1Observational Cosmology Lab., Code 665, NASA Goddard Space Flight Center, Greenbelt MD 20910, USA 2Department of Astronomy, University of Maryland, College Park, MD 20742, USA 3Service d’Astrophysique, L’Orme des Merisiers, CEA/Saclay, 91191 Gif-sur-Yvette, France 4NASA Ames Research Center, Mail Stop 245-3, Moffett Field, CA 94035, USA 5NASA Ames Research Center, Mail Stop 245-6, Moffett Field, CA 94035, USA 6SETI Intstitute, 515 N. Whisman Rd, Mountain View, CA 94043, USA 7Physics & Astronomy Dept., University of Western Ontario, PAB 213, London ON N6A 3K7, Canada 8Institut d’Astrophysique Spatiale, Université de Paris XI, 91405 Orsay, France

We present the results of a systematic study of mid-IR spectra of Galactic regions, Magellanic H ii regions, and galaxies of various types (dwarf, spiral, starburst), observed by the satellites ISO and Spitzer. We study the relative variations of the 6.2, 7.7, 8.6 and 11.3 μm features inside spatially resolved objects (such as M 82, M 51, 30 Dor, M 17 and the Orion bar, as well as among 90 integrated spectra of 50 objects. Our main results are that the 6.2, 7.7 and 8.6 μm bands are essentially tied together, while the ratios between these bands and the 11.3 μm band varies by one order of magnitude. This implies that the properties of the PAHs are remarkably universal throughout our sample, and that the relative variations of the band ratios are mainly controled by the fraction of ionized PAHs. In particular, we show that we can rule out both the modification of the PAH size distribution, and the mid-infrared extinction, as an explanation of these variations. Using a few well-studied Galactic regions (including the spectral image of the Orion bar), we give an empirical relation between the I_{6.2}/I_{11.3} ratio and the ionization/recombination ratio G_{0}/n_{e} ×\sqrt{T_{gas}}, therefore providing a useful quantitative diagnostic tool of the physical conditions in the regions where the PAH emission originates. Finally, we discuss the physical interpretation of the I_{6.2}/I_{11.3} ratio, on galactic size scales.

Accepted for publication in ApJ Available from arXiv:0801.4955

The Origin of the Magellanic Stream and Its Leading Arm

David L. Nidever^{1}, Steven R. Majewski^{1} and W. Butler Burton^{2,3}

1Department of Astronomy, University of Virginia, Charlottesville, VA, USA 2Sterrewacht Leiden, Leiden, The Netherlands 3National Radio Astronomy Observatory, Charlottesville, VA, USA

We explore the Magellanic Stream (MS) using a Gaussian decomposition of the H i velocity profiles in the Leiden-Argentine-Bonn (LAB) all-sky H i survey. This decomposition exposes the MS to be composed of two filaments distinct both spatially (as first pointed out by Putman et al.) and in velocity. Using the velocity coherence of the filaments, one can be traced back to its origin in what we identify as the SouthEast H i Overdensity (SEHO) of the Large Magellanic Cloud (LMC), which includes 30 Doradus. Parts of the Leading Arm (LA) can also be traced back to the SEHO in velocity and position. Therefore, at least one-half of the trailing Stream and most of the LA originates in the LMC, contrary to previous assertions that both the MS and the LA originate in the Small Magellanic Cloud (SMC) and/or in the Magellanic Bridge. The two MS filaments show strong periodic, undulating spatial and velocity patterns that we speculate are an imprint of the LMC rotation curve. If true, then the drift rate of the Stream gas away from the Magellanic Clouds is ∼49 km s^{−1} and the age of the MS is ∼1.74 Gyr. The Staveley-Smith et al. high-resolution H i data of the LMC show gas outflows from supergiant shells in the SEHO that seem to be creating the LA and LMC filament of the MS. Blowout of LMC gas is an effect not previously accounted for but one that probably plays an important role in creating the MS and LA.

Accepted for publication in The Astrophysical Journal Available from arXiv:0706.1578

The Iron Abundance of the Magellanic Bridge

P.L. Dufton^{1}, R.S.I. Ryans^{1}, H. Thompson^{1} and R. Street^{2}

1Queen’s University of Belfast, UK 2University of California, USA

High-resolution HST ultra-violet spectra for five B-type stars in the Magellanic Bridge and in the Large and Small Magellanic Clouds have been analysed to estimate their iron abundances. Those for the Clouds are lower than estimates obtained from late-type stars or the optical lines in B-type stars by approximately 0.5 dex. This may be due to systematic errors possibly arising from non-LTE effects or from errors in the atomic data as similar low Fe abundances having previously been reported from the analysis of the ultra-violet spectra of Galactic early-type stars.

The iron abundance estimates for all three Bridge targets appear to be significantly lower than those found for the SMC and LMC by approximately −0.5 dex and −0.8 dex respectively and these differential results should not be affected by any systematic errors present in the absolute abundance estimates. These differential iron abundance estimates are consistent with the underabundances for C, N, O, Mg and Si of approximately −1.1 dex relative to our Galaxy previously found in our Bridge targets. The implications of these very low metal abundances for the Magellanic Bridge are discussed in terms of metal deficient material being stripped from the SMC.

Accepted for publication in MNRAS Available from arXiv:0802.0132

Little Dwarf Galaxies survive within Bigger Dwarfs: Why Some Dwarfs Go Dark and Others Stay Luminous

George Lake^{1} and Elena D’Onghia^{1}

1University of Zurich, Switzerland

The Large Magellanic Cloud (LMC) was the largest galaxy in a group of galaxies that entered the Milky Way (MW) system at late times. Seven of the 11 brightest dwarf galaxies of the MW may have been part of this system. The association of dwarfs with the plane of the LMCs orbit has been used to argue that they are formed from tidal debris from LMC and SMC (Kroupa et al. 2005). Instead, we find that they owe to the tidal breakup of the Magellanic Group. The Cold Dark Matter (CDM) paradigm suffers from the small-scale structure problem where 500 galaxies as massive as Draco and Ursa Minor are expected, but only 11 are seen. If seven of the 11 observed were part of the LMC group, the substructure in this group is close to CDM predictions. There are other likely nearby dwarf groups, including a second Milky Way group associated with Fornax (Lynden-Bell 1982), great circles of satellites in M31 (Koch & Grebel 2006) and groupings of otherwise isolated dwarfs (Tully et al. 2006). For the first time, we build he circular velocity distribution of the satellites in system as small as groups of dwarf galaxies, using the dwarf galaxies that likely entered with the LMC system and data on nearby dwarf systems. Our work points to natural mechanisms that lead to less suppression of satellites in dwarf groups providing an explanation for the missing satellite problem in the Local Group.

Submitted to ApJ Letters Available from arXiv:0802.0001

The Persistence of Memory, or How the X-Ray Spectrum of SNR 0509−67.5 Reveals the Brightness of its Parent Type Ia Supernova

Carles Badenes^{1,2}, John P. Hughes^{2}, Gamil Cassam-Chenai^{2} and Eduardo Barvo^{3}

1Princeton University, USA 2Rutgers University, USA 3Universitat Politécnica de Catalunya/Institut d’Estudis Espacials de Catalunya, Spain

We examine the dynamics and X-ray spectrum of the young Type Ia supernova remnant 0509−67.5 in the context of the recent results obtained from the optical spectroscopy of its light echo. Our goal is to estimate the kinetic energy of the supernova explosion using Chandra and XMM-Newton observations of the supernova remnant, thus placing the birth event of 0509−67.5 in the sequence of dim to bright Type Ia supernovae. We base our analysis on a standard grid of one-dimensional delayed detonation explosion models, together with hydrodynamic and X-ray spectral calculations of the supernova remnant evolution. From the remnant dynamics and the properties of the O, Si, S, and Fe emission in its X-ray spectrum we conclude that 0509−67.5 was originated ∼400 years ago by a bright, highly energetic Type Ia explosion similar to SN 1991T. Our best model has a kinetic energy of 1.4 × 10^{51} erg and synthesizes 0.97 M_{⊙} of 56Ni. These results are in excellent agreement with the age estimate and spectroscopy from the light echo. We have thus established the first connection between a Type Ia supernova and its supernova remnant based on a detailed quantitative analysis of both objects.

Accepted for publication in ApJ Available from arXiv:0801.4761

An X-ray Survey of Wolf-Rayet Stars in the Magellanic Clouds. I. The Chandra ACIS Dataset

Martín A. Guerrero^{1,2} and You-Hua Chu^{2}

1Instituto de Astrofísica de Andalucía, IAA-CSIC, Spain 2University of Illinois at Urbana-Champaign, UIUC, USA

Wolf-Rayet (WR) stars are evolved massive stars with strong fast stellar winds. WR stars in our Galaxy have shown three possible sources of X-ray emission associated with their winds: shocks in the winds, colliding stellar winds, and wind-blown bubbles; however, quantitative analyses of observations are often hampered by uncertainties in distances and heavy foreground absorption. These problems are mitigated in the Magellanic Clouds (MCs), which are at known distances and have small foreground and internal extinction. We have therefore started a survey of X-ray emission associated with WR stars in the MCs using archival Chandra, ROSAT, and XMM-Newton observations. In the first paper of this series, we report the results for 70 WR stars in the MCs using 192 archival Chandra ACIS observations. X-ray emission is detected from 29 WR stars. We have investigated their X-ray spectral properties, luminosities, and temporal variability. These X-ray sources all have luminosities greater than a few times 10^{32} ergs s^{−1}, with spectra indicative of highly absorbed emission from a thin plasma at high temperatures typical of colliding winds in WR+OB binary systems. Significant X-ray variability with periods ranging from a few hours up to ∼20 days is seen associated with several WR stars. In most of these cases, the X-ray variability can be linked to the orbital motion of the WR star in a binary system, further supporting the colliding wind scenario for the origin of the X-ray emission from these stars.

Accepted for publication in The Astrophysical Journal Supplement Series Available from arXiv:0802.0503

An X-ray Survey of Wolf-Rayet Stars in the Magellanic Clouds. II. The ROSAT PSPC and HRI Datasets

Martín A. Guerrero^{1,2} and You-Hua Chu^{2}

1Instituto de Astrofísica de Andalucía, IAA-CSIC, Spain 2University of Illinois at Urbana-Champaign, UIUC, USA

Wolf-Rayet (WR) stars in the Magellanic Clouds (MCs) are ideal for studying the production of X-ray emission by their strong fast stellar winds. We have started a systematic survey for X-ray emission from WR stars in the MCs using archival Chandra, ROSAT, and XMM-Newton observations. In Paper I, we reported the detection of X-ray emission from 29 WR stars using Chandra ACIS observations of 70 WR stars in the MCs. In this paper, we report the search and analysis of archival ROSAT PSPC and HRI observations of WR stars. While useful ROSAT observations are available for 117 WR stars in the MCs, X-ray emission is detected from only 7 of them. The detection rate of X-ray emission from MCs WR stars in the ROSAT survey is much smaller than in the Chandra ACIS survey, illustrating the necessity of high angular resolution and sensitivity. LMC-WR 101-102 and 116 were detected by both ROSAT and Chandra, but no large long-term variations are evident.

Accepted for publication in The Astrophysical Journal Supplement Series Available from arXiv:0802.0493

Black holes and core expansion in massive star clusters

A.D. Mackey^{1}, M.I. Wilkinson^{2}, M.B. Davies^{3} and G.F. Gilmore^{4}

1University of Edinburgh, Scotland 2University of Leicester, England 3Lund University, Sweden 4University of Cambridge, England

We present the results from realistic N-body modelling of massive star clusters in the Magellanic Clouds. We have computed eight simulations with N ∼ 10^{5} particles; six of these were evolved for at least a Hubble time. The aim of this modelling is to examine the possibility of large-scale core expansion in massive star clusters and search for a viable dynamical origin for the radius-age trend observed for such objects in the Magellanic Clouds. We identify two physical processes which can lead to significant and prolonged cluster core expansion: mass-loss due to rapid stellar evolution in a primordially mass segregated cluster, and heating due to a retained population of stellar-mass black holes. These two processes operate over different time-scales — the former occurs only at early times and cannot drive core expansion for longer than a few hundred Myr, while the latter typically does not begin until several hundred Myr have passed but can result in core expansion lasting for many Gyr. We investigate the behaviour of these expansion mechanisms in clusters with varying degrees of primordial mass segregation and in clusters with varying black hole retention fractions. In combination, the two processes can lead to a wide variety of evolutionary paths on the radiusage plane, which fully cover the observed cluster distribution and hence define a dynamical origin for the radius-age trend in the Magellanic Clouds. We discuss the implications of core expansion for various aspects of globular cluster research, as well as the possibility of observationally inferring the presence of a population of stellar-mass black holes in a cluster.

Accepted for publication in MNRAS Available from arXiv:0802.0513

Breaking up the Magellanic Group into the Milky Way Halo: Understanding the Local Dwarf Galaxy Properties

Elena D’Onghia^{1}

1University of Zurich, Switzerland

We use a numerical simulation of a loose group containing a Milky Way halo to probe that in the hierarchical universe the Magellanic Clouds and some dSphs have been accreted into the Milky Way halo from a late infalling group of dwarfs. Our simulations show that the tidal breakup of the Magellanic group occurs before it enters the Milky Way halo. Only half of the satellites contributed from the group are predicted to be inside the Milky Way virial radius. Half of its subhalos survive outside the current virial radius in the form of satellites, whereas the remaining material contributes to the diffuse Milky Way halo. At z ∼ 0 the disrupted group contributes less than 10% to the Milky Way halo mass but 20% of the brightest dwarf galaxies of the Milky Way have been part of this group. This scenario points out that some dSphs might have been formed away from giant spirals and been accreted already as spheroids, by a late infall group in contrast with the classical picture of tidal stripping of dSph formation models. This would naturally explain several peculiarities of the local dSph: why Draco and the other luminous dSphs exist compared to other ultra-faint satellite galaxies, the location of Tucana and Cetus in the outskirts of the Local Group and the mismatch in metallicity between the stellar halo of the Milky Way and the dwarf galaxies that many have suspected dissolved to build it.

Submitted to ApJ Letters Available from arXiv:0802.0302

Submillimeter Line Emission from LMC N159W: a dense, clumpy PDR in a low metallicity environment

J. L. Pineda^{1}, N. Mizuno^{2}, J. Stutzki^{3}, M. Cubick^{3}, M. Aravena^{1}, F. Bensch^{1}, F. Bertoldi^{1}, L. Bronfman^{4}, K.

Fujishita^{2}, U.U. Graf^{3}, M. Hitschfeld^{3}, N. Honingh^{3}, H. Jakob^{3}, K. Jacobs^{3}, A. Kawamura^{2}, U. Klein^{1}, C. Kramer^{3}, J. May^{4}, M. Miller^{3}, Y. Mizuno^{2}, P. Müller^{1}, T. Onishi^{2}, V. Ossenkopf^{3}, D. Rabanus^{3}, M. Röllig^{3}, M.

Rubio^{4}, H. Sasago^{2}, R. Schieder^{3}, R. Simon^{3}, K. Sun^{3}, N. Volgenau^{3} and Y. Fukui^{2}

1Argelander-Institut für Astronomie, Auf dem Hügel 71, D-53121 Bonn, Germany 2Department of Astrophysics, Nagoya University, Chikusa-ku, Nagoya 464-8602, Japan 3KOSMA, I. Physikalisches Institut, Universität zu Köln, Zülpicher Straße 77, D-50937 Köln, Germany 4Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile

Star formation at earlier cosmological times took place in an interstellar medium with low metallicity. The Large Magellanic Cloud (LMC) is ideally suited to study star formation in such an environment. The physical and chemical state of the ISM in a star forming environment can be constrained by observations of submm and FIR spectral lines of the main carbon carrying species, CO, C i and C ii, which originate in the surface layers of molecular clouds illuminated by the UV radiation of the newly formed, young stars. We present high-angular resolution sub-millimeter observations in the N159W region in the LMC obtained with the NANTEN2 telescope of the ^{12}CO J = 4 → 3, J = 7 → 6, and 13CO J = 4 → 3 rotational and [C i] ^{3}P_{1}−^{3}P_{0} and ^{3}P_{2}−^{3}P_{1} fine-structure transitions. The ^{13}CO J = 4 → 3 and [C i] ^{3}P_{2}−^{3}P_{1} transitions are detected for the first time in the LMC. We derive the physical and chemical properties of the low-metallicity molecular gas using an escape probability code and a self-consistent solution of the chemistry and thermal balance of the gas in the framework of a clumpy cloud PDR model. The separate excitation analysis of the submm CO lines and the carbon fine structure lines shows that the emitting gas in the N159W region has temperatures of about 80 K and densities of about 10^{4} cm^{−3}. The estimated C to CO abundance ratio close to unity is substantially higher than in dense massive star-forming regions in the Milky Way. The analysis of all observed lines together, including the [C ii] line intensity reported in the literature, in the context of a clumpy cloud PDR model constrains the UV intensity to about χ ≈ 220 and an average density of the clump ensemble of about 10^{5} cm^{−3}, thus confirming the presence of high density material in the LMC N159W region.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:0802.1929 and from http://www.astro.uni-bonn.de/∼jopineda/pega/n159w paper.pdf

The many streams of the Magellanic Stream

Snežana Stanimirović^{1}, Samantha Hoffman^{1}, Carl Heiles^{2}, Kevin A. Douglas^{3}, Mary Putman^{4} and Joshua E. G. Peek^{2}

1Department of Astronomy, University of Wisconsin, Madison, WI 53706, USA 2Department of Astronomy, UC Berkeley, 601 Campbell Hall, Berkeley, CA 9472, USA 3Space Sciences Laboratory, University of California, Berkeley, CA 94720, USA 4University of Michigan, Department of Astronomy, 500 Church St., Ann Arbor, MI 48109, USA

We present results from neutral hydrogen (H i) observations of the tip of the Magellanic Stream (MS), obtained with the Arecibo telescope as a part of the on-going survey by the Consortium for Galactic studies with the Arecibo L-band Feed Array. We find four large-scale, coherent H i streams, extending continuously over a length of 20^{∘}, each stream possessing different morphology and velocity gradients. The newly discovered streams provide strong support for the tidal model of the MS formation by Connors et al. (2006), which suggested a spatial and kinematic bifurcation of the MS. The observed morphology and kinematics suggest that three of these streams could be interpreted as a 3-way splitting of the main MS filament, while the fourth stream appears much younger and may have originated from the Magellanic Bridge. We find an extensive population of H i clouds at the tip of the MS. Two thirds of clouds have an angular size in the range 3.5′–10′. We interpret this as being due to thermal instability, which would affect a warm tail of gas trailing through the Galactic halo over a characteristic timescale of a few Myrs to a few hundred Myrs. We show that thermal fragments can survive in the hot halo for a long time, especially if surrounded by a < 10^{6} K halo gas. If the observed clumpy structure is mainly due to thermal instability, then the tip of the MS is at a distance of ∼ 70 kpc. A significant fraction of H i clouds at the tip of the MS show multi-phase velocity profiles, indicating the co-existence of cooler and warmer gas.

Accepted for publication in Astrophysical Journal Available from arXiv:0802.1349

The VLT-FLAMES survey of massive stars: rotation and nitrogen enrichment as the key to understanding massive star evolution

I. Hunter^{1,2}, I. Brott^{3}, D.J. Lennon^{4}, N. Langer^{3}, P.L. Dufton^{1}, C. Trundle^{1}, S.J. Smartt^{1}, A. de Koter^{5}, C.J. Evans^{6} and R.S.I. Ryans^{1}

1Astrophysical Research Centre, School of Mathematics and Physics, Queen’s Univeristy of Belfast, Belfast, BT7 1NN, N. Ireland 2Isaac Newton Group of Telescopes, Apartado de Correos 321, 38700 Santa Cruz de La Palma, Canary Islands, Spain 3Astronomical Institute, Utrecht University, Princetonplein 5, NL-3584CC, Utrecht, The Netherlands 4Instituto de Astrofísica de Canarias, Calle Vía Láctea, E-38200 La Laguna, Spain 5Astronomical Institute Anton Pannekoek, University of Amsterdam, Kruislaan 403, 1098 SJ Amsterdam, The Netherlands 6UK Astronomy Technology Centre, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, Scotland

Rotation has become an important element in evolutionary models of massive stars, specifically via the prediction of rotational mixing. Here, we study a sample of stars, including rapid rotators, to constrain such models and use nitrogen enrichments as a probe of the mixing process. Chemical compositions (C, N, O, Mg and Si) have been estimated for 135 early B-type stars in the Large Magellanic Cloud with projected rotational velocities up to ∼300 km s^{−1} using a non-LTE tlusty model atmosphere grid. Evolutionary models, including rotational mixing, have been generated attempting to reproduce these observations by adjusting the overshooting and rotational mixing parameters and produce reasonable agreement with 60% of our core hydrogen burning sample. We find (excluding known binaries) a significant population of highly nitrogen enriched intrinsic slow rotators (v sin i< ∼50 km s^{−1}) incompatible with our models (∼20% of the sample). Furthermore, while we find fast rotators with enrichments in agreement with the models, the observation of evolved (log g < 3.7 dex) fast rotators that are relatively unenriched (a further ∼20% of the sample) challenges the concept of rotational mixing. We also find that 70% of our blue supergiant sample cannot have evolved directly from the hydrogen burning main-sequence. We are left with a picture where invoking binarity and perhaps fossil magnetic fields are required to understand the surface properties of a population of massive main sequence stars.

Accepted for publication in ApJ Letters Available from arXiv:0711.2267

The VLT-FLAMES survey of massive stars: atmospheric parameters and rotational velocity distributions for B-type stars in the Magellanic

Clouds

I. Hunter^{1,2}, D.J. Lennon^{2,3}, P.L. Dufton^{1}, C. Trundle^{1}, S. Simón-Díaz^{4}, S.J. Smartt^{1}, R.S.I. Ryans^{1} and C.J. Evans^{5}

1Astrophysical Research Centre, School of Mathematics and Physics, Queen’s Univeristy of Belfast, Belfast, BT7 1NN, N. Ireland 2Isaac Newton Group of Telescopes, Apartado de Correos 321, 38700 Santa Cruz de La Palma, Canary Islands, Spain 3Instituto de Astrofísica de Canarias, 38200 La Laguna, Tenerife, Spain 4LUTH, Observatoire de Meudon, 5 Place Jules Janssen, 92195, Meudon Cedex, France 5UK Astronomy Technology Centre, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, Scotland

Aims: We aim to provide the atmospheric parameters and rotational velocities for a large sample of O- and early B-type stars, analysed in a homogeneous and consistent manner, for use in constraining theoretical models. Methods: Atmospheric parameters, stellar masses, and rotational velocities have been estimated for approximately 250 early B-type stars in the Large (LMC) and Small (SMC) Magellanic Clouds from high-resolution VLT-FLAMES data using the non-LTE TLUSTY model atmosphere code. This data set has been supplemented with our previous analyses of some 50 O-type stars (Mokiem et al. 2006, 2007) and 100 narrow-lined early B-type stars (Hunter et al. 2006, Trundle et al. 2007) from the same survey, providing a sample of ∼400 early-type objects. Results: Comparison of the rotational velocities with evolutionary tracks suggests that the end of core hydrogen burning occurs later than currently predicted and we argue for an extension of the evolutionary tracks. We also show that the large number of the luminous blue supergiants observed in the fields are unlikely to have directly evolved from main-sequence massive O-type stars as neither their low rotational velocities nor their position on the H-R diagram are predicted. We suggest that blue loops or mass-transfer binary systems may populate the blue supergiant regime. By comparing the rotational velocity distributions of the Magellanic Cloud stars to a similar Galactic sample, we find that (at 3σ confidence level) massive stars (above 8 M_{⊙}) in the SMC rotate faster than those in the solar neighbourhood. However there appears to be no significant difference between the rotational velocity distributions in the Galaxy and the LMC. We find that the v sin i distributions in the SMC and LMC can modelled with an intrinsic rotational velocity distribution that is a Gaussian peaking at 175 km s^{−1} (SMC) and 100 km s^{−1} (LMC) with a \frac{1}{e} half width of 150 km s^{−1}. We find that in NGC 346 in the SMC, the 10–25 M_{⊙} main-sequence stars appear to rotate faster than their higher mass counterparts. It is not expected that O-type stars spin down significantly through angular momentum loss via stellar winds at SMC metallicity, hence this could be a reflection of mass dependent birth spin rates. Recently Yoon et al. (2006) have determined rates of GRBs by modelling rapidly rotating massive star progenitors. Our measured rotational velocity distribution for the 10–25 M_{⊙} stars is peaked at slightly higher velocities than they assume, supporting the idea that GRBs could come from rapid rotators with initial masses as low as 14 M_{⊙} at low metallicities.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:0711.2264

Chemical composition and mixing in giant H ii regions: NGC 3603, 30 Doradus, and N 66

V. Lebouteiller^{1}, J. Bernard-Salas^{1}, B. Brandl^{2}, D.G. Whelan^{1}, Yanling Wu^{1}, V. Charmandaris^{3}, D. Devost^{1} and J.R. Houck^{1}

1Cornell University, USA 2Leiden Observatory, The Netherlands 3Department of Physics, University of Crete, Greece

We investigate the chemical abundances of NGC 3603 in the Milky Way, of 30 Doradus in the Large Magellanic Cloud, and of N 66 in the Small Magellanic Cloud. Mid-infrared observations with the Infrared Spectrograph onboard the Spitzer Space Telescope allow us to probe the properties of distinct physical regions within each object: the central ionizing cluster, the surrounding ionized gas, photodissociation regions, and buried stellar clusters. We detect [S iii], [S iv], [Ar iii], [Ne ii], [Ne iii], [Fe ii], and [Fe iii] lines and derive the ionic abundances. Based on the ionic abundance ratio (Ne iii/H)/(S iii/H), we find that the gas observed in the MIR is characterized by a higher degree of ionization than the gas observed in the optical spectra. We compute the elemental abundances of Ne, S, Ar, and Fe. We find that the alpha-elements Ne, S, and Ar scale with each other. Our determinations agree well with the abundances derived from the optical. The Ne/S ratio is higher than the solar value in the three giant H ii regions and points toward a moderate depletion of sulfur on dust grains. We find that the neon and sulfur abundances display a remarkably small dispersion (0.11 dex in 15 positions in 30 Doradus), suggesting a relatively homogeneous ISM, even though small-scale mixing cannot be ruled out.

Accepted for publication in ApJ Available from arXiv:0710.4549

Evidence for an Overluminosity of the Variable Star RR Lyr, and a Revised Distance to the LMC

M. Catelan^{1} and C. Cortés^{1,2}

1Pontificia Universidad Católica de Chile, Santiago, Chile 2Universidade Federal do Rio Grande do Norte, Natal, Brazil

We use theoretical models to establish a tight relationship for the absolute magnitudes of RR Lyrae stars as a function of their periods and Strömgren pseudo-color c_{0} ≡ (u−v)_{0} − (v −b)_{0}. Applying this to RR Lyr, and comparing the result with the predicted average absolute magnitude for stars of similar metallicity from the same models, yields an overluminosity of 0.064 ± 0.013 mag in Strömgren y (and thus similarly in V) for RR Lyr. Based on a revised value for RR Lyr’s trigonometric parallax, and on a newly derived reddening value of E(B − V) = 0.015 ± 0.020, we provide a corrected relationship between average absolute magnitude and metallicity for RR Lyrae stars that takes RR Lyr’s evolutionary status fully into account for the first time. Applying this relationship to the LMC, we derive a revised true distance modulus of (m − M)_{0} = 18.44 ± 0.11.

Accepted for publication in The Astrophysical Journal Letters Available from arXiv:0802.2063

A Long Look at the Be/X-Ray Binaries of the Small Magellanic Cloud

J.L. Galache^{1}, R.H.D. Corbet^{2,3}, M.J. Coe^{4}, S. Laycock^{1}, M.P.E. Schurch^{4}, C. Markwardt^{2,5}, F.E. Marshall^{2} and J. Lochner^{2,6}

1Harvard-Smithsonian Center for Astrophysics, USA 2NASA Goddard Space Flight Center, USA 3University of Maryland, Baltimore County, USA 4University of Southampton, UK 5University of Maryland, College Park, USA 6Universities Space Research Association

We have monitored 41 BeX binary systems in the Small Magellanic Cloud over ∼9 years using PCA-RXTE data from a weekly survey program. The resulting light curves were analysed in search of orbital modulations with the result that 10 known orbital ephemerides were confirmed and refined, while 10 new ones where determined. A large number of X-ray orbital profiles are presented for the first time, showing similar characteristics over a wide range of orbital periods. Lastly, three pulsars: SXP 46.4, SXP 89.0 and SXP 165 were found to be misidentifications of SXP 46.6, SXP 91.1 and SXP 169, respectively.

Accepted for publication in ApJS Available from arXiv:0802.2118

Chandra HETG Spectra of SN 1987A at 20 Years

D. Dewey^{1}, S.A. Zhekov^{2,3}, R. McCray^{2} and C.R. Canizares^{1}

1MIT Kavli Institute, USA 2JILA, University of Colorado, USA 3On leave from Space Research Institute, Sofia, Bulgaria

We have undertaken deep, high-resolution observations of SN 1987A at ∼20 years after its explosion with the Chandra HETG and LETG spectrometers. Here we present the HETG X-ray spectra of SN 1987A having unprecedented spectral resolution and signal-to-noise ratios in the 6–20 Å bandpass, which includes the H-like and He-like lines of Si, Mg, and Ne, as well as O VIII lines and bright Fe xvii lines. In joint analysis with LETG data, we find that there has been a significant decrease from 2004 to 2007 in the average temperature of the highest temperature component of the shocked-plasma emission. Model fitting of the profiles of individual HETG lines yields bulk kinematic velocities of the higher-Z ions, Mg and Si, that are significantly lower than those inferred from the LETG 2004 observations.

Accepted for publication in ApJ, 676, L131 (2008) Available from arXiv:0802.2340 and from http://space.mit.edu/home/dd

Milky Way potentials in CDM and MOND. Is the Large Magellanic Cloud on a bound orbit?

Xufen Wu^{1}, Benoit Famaey^{2}, Gianfranco Gentile^{3,4}, Hagai Perets^{5} and HongSheng Zhao^{1}

1SUPA, School of Physics and Astronomy, University of St. Andrews, Fife KY16 9SS, UK 2Institut d’Astronomie et d’Astrophysique, Université Libre de Bruxelles,CP 226, Boulevard du Triomphe, B-1050, Bruxelles, Belgium 3University of New Mexico, Department of Physics and Astronomy, 800 Yale Boulevard NE, Albuquerque, NM 87131, USA 4Sterrenkundig Observatorium, Ghent University, Krijgslaan 281, S9, B-9000 Ghent, Belgium 5Benoziyo centre for Astrophysics, Weizmann Institute of Science, POB 26, Rehovot, Israel

We compute the Milky Way potential in different cold dark matter (CDM) based models, and compare these with the modified Newtonian dynamics (MOND) framework. We calculate the axis ratio of the potential in various models, and find that isopotentials are less spherical in MOND than in CDM potentials. As an application of these models, we predict the escape velocity as a function of the position in the Galaxy. This could be useful in comparing with future data from planned or already-underway kinematic surveys (RAVE, SDSS, SEGUE, SIM, GAIA or the hypervelocity stars survey). In addition, the predicted escape velocity is compared with the recently measured high proper motion velocity of the Large Magellanic Cloud (LMC). To bind the LMC to the Galaxy in a MOND model, while still being compatible with the RAVE-measured local escape speed at the Sun’s position, we show that an external field modulus of less than 0.03a_{0} is needed.

Accepted for publication in MNRAS Available from arXiv:0803.0977

The Luminosities and Distance Scales of Type II Cepheid and RR Lyrae Variables

Michael W. Feast^{1}, Clifton D. Laney^{2}, Thomas D. Kinman^{3}, Floor van Leeuwen^{4} and Patricia A. Whitelock^{1,2,5}

1Astronomy Dept University of Cape Town, South Africa 2South African Astronomical Observatory, Cape Town, South Africa 3National Optical Observatory, Tuscon, U.S.A. 4Institute of Astronomy, Cambridge, U.K. 5National Astrophysics and Space Science Programme, University of Cape Town, South Africa

Infrared and optical absolute magnitudes are derived for the Type II Cepheids κ Pav and VY Pyx using revised Hipparcos parallaxes and for κ Pav, V553 Cen and SW Tau from pulsation parallaxes. Revised Hipparcos and HST (Benedict et al.) parallaxes agree satisfactorily and are combined in deriving absolute magnitudes. Phase-corrected J,H,K_{s} mags are given for 142 Hipparcos RR Lyraes based on 2MASS observations. Pulsation and trigonometrical parallaxes for classical Cepheids are compared to establish the best value for the projection factor (p) used in pulsational analyses. The Mv of RR Lyrae itself is 0.16 ± 0.12 mag brighter than predicted from a M_{V}-[Fe/H] relation based on RR Lyrae stars in the LMC at a modulus of 18.39 ± 0.05 as found from classical Cepheids. This is consistent with the prediction of Catelan and Cortes that it is overluminous for its metallicity. The M(K) results for the metal and carbon rich, Galactic disc stars, V553 Cen and SW Tau, each with a small internal error (±0.08 mag), have a mean deviation of only 0.02 mag from the Period-Luminosity relation established by Matsunaga et al. for Type II Cepheids in globular clusters and with a zero-point based on the same LMC scale. Comparing directly the luminosities of these two stars with published data on Type II Cepheids in the LMC and in the Galactic Bulge leads to an LMC modulus of 18.37 ± 0.09 and a distance to the Galactic Centre of R_{o} = 7.64 ± 0.21 kpc. The data for VY Pyx agrees with these results within the uncertainties set by its parallax. Evidence is presented that κ Pav may have a close companion and possible implications of this are discussed. If the pulsation parallax of this star is incorporated in the analyses, the distance scales just discussed will be increased by about 0.15 ± 0.15 mag. V553 Cen and SW Tau show that at optical wavelengths PL relations are wider for field stars than for those in globular clusters. This is probably due to a narrower range of masses in the latter case.

Accepted for publication in MNRAS Available from arXiv:0803.0466

Conference Paper

On the Cepheid Metallicity Dichotomy

Geza Kovacs^{1}

1Konkoly Observatory, Budapest, Hungary

With the aid of stellar atmosphere models, we derive iron abundances [Fe/H] from the OGLE B, V, I photometry on the Magellanic Cloud Cepheids. We show that in both clouds the average metallicities of the first overtone variables are lower than those of the fundamental ones (by ∼0.2 and ∼0.3 dex in the LMC and SMC, respectively). Consequently, there is a correlation between the overall [Fe/H] and luminosity; the lower luminosity stars tend to be also of lower metallicity. These metallicity dependencies are concordant with the ones derived for the two types of double-mode Cepheids from pulsation theory. Yet another support of this dichotomy comes from the evolution theories that require lower metallicities for blue-looping low-luminosity stars than for high-luminosity ones. We also comment on the possibility of using period-luminosity-color relations to derive more accurate metallicities.

Oral contribution, published in a workshop held at IAP, Paris, July 9–11, 2008: ”Nonlinear stellar hydrodynamics and the pulsations of Cepheids”, Eds. Goupil M-J., Kollath Z., Kervella P., & Nardetto N., Publ.: EDP Sciences Available from arXiv:0802.4166

Job Advert

University of Manchester Research Associate in Astrophysics (Post-doc)

The Jodrell Bank Centre for Astrophysics, based in the Alan Turing Building at the University of Manchester, invites applications for a postdoctoral research associate to work on the SAGE-Spec project funded by STFC. The successful candidate will be expected to manage the content of the spectral database, and will contribute to the spectral classification. The postholder will also pursue research using the SAGE-Spec data to investigate the properties of the life cycle of dust and gas in the Large Magellanic Cloud.

See also http://www.manchester.ac.uk/aboutus/jobs/research/vacancy/index.htm?ref=132996